Jeff, Tom, and Chelsea order a large pizza. They decide to split the cost according to how much they will eat. Tom pays dollar-sign 5.49, Chelsea eats StartFraction 1 Over 8 EndFraction of the pizza, and Jeff eats StartFraction 1 Over 2 EndFractionof the pizza. How much did the pizza cost?
Answer by macston(5194)

.
Chelsea eats 1/8 and Jeff eats 1/2
1-(1/8)-(1/2)=3/8
Tom ate 3/8.
($5.49)/(3/8)=$14.64
ANSWER: Total cost of the pizza was $14.64.